# Measuring Agents in Production

### v1

@@ 4.2 Application Domains
We find that agents operate in diverse industries well beyond software engineering. Figure 2 shows that the three domains with the highest number of reported deployments in survey are Finance & Banking (39.1%), Technology (24.6%), and Corporate Services (23.2%). We also observe a long tail of applications in Data Analytics (13.0%), Research & Development (11.6%), and other specialized domains (15.9%). This distribution demonstrates that agentic systems deliver practical value across fundamentally different industries.
@@ 4.3 Users of Agents
We find that deployed agents primarily serve human users directly. Survey data indicates that 92.5% of deployed systems serve humans rather than other agents or software systems. As shown in Figure 5(a), internal employees are the primary user base (52.2%), followed by external customers (40.3%). Only 7.5% of deployed systems serve non-human consumers.
The focus on internal users is a deliberate deployment choice. Detailed case studies reveal that organizations restrict deployments to internal environments to mitigate unsolved reliability and security concerns. Internal users operate within organizational boundaries where agent mistakes have lower consequences and human oversight is readily available. An example from our case study, internal PagerDuty agents respond to employee requests, while human engineers can take over when needed.
Furthermore, most systems—including external-facing applications—require domain-specific knowledge to operate. For example, insurance authorization agents support nurses, and incident response agents support site reliability engineers. This reflects a pattern where agents function as tools that augment domain experts rather than replace them. This paradigm also enables human users to serve as the final verifiers of agent outputs, which we discuss further in Section 6.2.
Beyond user type, we examine the scale of the user base. We find that end-user counts for deployed systems vary significantly. As shown in Figure 4(c), 42.9% of deployments serve user bases in the hundreds. However, we also observe high-traffic deployments (25.7%) serving tens of thousands to over 1 million users daily, representing substantial user impact or possibly mature systems.
@@ 4.4 Latency Requirements
Relaxed latency requirements are common among deployed agents. Figure 5(b) shows the distribution of maximum allowable end-to-end latency. Minutes is the most common target, followed by seconds. Notably, 17.0% report no defined limit yet, and 1.9% allow hours to days.
The latency tolerance reflects the productivity focus from Section 4.1. Agents are often used to automate tasks that typically take humans hours to complete. Consequently, an agent taking multiple minutes to respond remains orders of magnitude faster than non-agentic baselines. Interview participants emphasized this advantage: even if an agent takes five minutes, that remains more than faster than assigning the task to a person on the team, especially when staffing shortages exist and the task is secondary to the user’s core responsibilities. Examples include nurses examining insurance details and software engineers responding to internal pager duty. Some deployed agents from case studies even batch requests hourly or overnight, further indicating latency is not a primary constraint.
However, this pattern breaks for real-time interactive applications. For example, practitioners building voice-driven systems report latency as their top challenge (Section 7.2) during detailed case study. These systems compete against human conversation speeds rather than task completion baselines. Among our 20 detailed case studies, only 5 require real-time responsiveness. The remaining 15 cases tolerate extended processing times: 7 involve human review with relaxed timing, 5 operate as asynchronous background processes, and 3 have hybrid operation patterns. For these systems, processing times of minutes remain acceptable because the alternative is days of human effort.
5 2 Results: What Models, Architectures, And Techniques Are Used To Build Agents?
Having established what problems practitioners target with agentic systems, we now address how these systems are built. We examine five critical implementation decisions: model selection, model weights tuning, prompt construction, agent architectures, and development frameworks.
Overall, practitioners favor established, straightforward methods over stochastic or training-intensive techniques. We find that methods popular in research—such as fine-tuning, reinforcement learning, and automated prompt optimization—remain uncommon in deployment. Instead, teams prioritize control, maintainability, and iteration speed.
@@ 5.1 Model Selection
We find that deployed agents rely heavily on proprietary models. Only 3 of 20 detailed case studies use open-source models (LLaMA 3 [grattafiori2024llama3herdmodels], Qwen [qwen3technicalreport], and GPT-OSS [openai2025gptoss]), while the remaining 17 rely on closed-source frontier models (Figure 6). Of the teams using closed-source models, 10 explicitly confirm using the Anthropic Claude [AnthropicClaude_nd] or OpenAI GPT [OpenAIclosed2025] families. Specific models mentioned during interviews include Anthropic Claude Sonnet 4 and Opus 4.1, and OpenAI o3, each representing the state-of-the-art model from each provider at the time of the interview. While other participants did not disclose specific models, most interviewees followed the pattern of using the largest and most capable state-of-the-art models as the primary model for their agents from each model family.
We find that open-source adoption is rare and is driven by specific constraints rather than general preference. Among the three cases using open-source models, motivations include high-volume workloads where inference costs at scale are prohibitive, and regulatory requirements preventing data sharing with external providers. For example, one team from detailed case study serve continuous infrastructure maintenance agent leverages the company’s existing internal compute resources (e.g., GPUs) to serve a fine-tuned open-source model to meet cost constraints.
For the majority of cases, model selection follows a pragmatic, empirical approach focused on downstream performance. Interviewees report that they test the top accessible state-of-the-art models for each task and select based on performance. Unlike the high-volume open-source use cases, these teams note that runtime costs are negligible compared to the human experts (e.g., medical professionals, senior engineers) that the agent augments. Consequently, they default to the best-performing closed-source models regardless of inference cost.†††Rankings based on public model leaderboards at the time of development. Additionally, 4 out of 20 detailed case studies combine LLMs with specialized off-the-shelf models (e.g., text-to-speech, chemistry foundation models) to handle specific modalities.
Number of Distinct Models. While a substantial portion rely on a single model, the majority coordinate multiple models to meet functional or operational needs. Survey results show that 40.9% of deployed agents use exactly one model, while 27.3% use two, 18.2% use three, and 13.6% use four or more. Among detailed case studies, 10 out of 20 (50%) combine models to address specific functional needs. We identify two drivers: cost optimization and modality. First, teams combine models of varying sizes to balance latency, cost, and quality. For example, one agent workflow from case study routes simple subtasks like pattern recognition to smaller models while reserving larger models for subtasks requiring higher reasoning capabilities. Second, teams integrate models to handle distinct data modalities. Communication agents leverage text-to-speech models alongside LLMs, while scientific agents employ domain-specific foundation models (e.g., chemistry) alongside general-purpose reasoning models.
Operational constraints also drive multi-model adoption. We find that multi-model architectures can emerge from lifecycle management needs rather than complex reasoning requirements for the agent task. Detailed case studies reveal that teams maintain multiple models to manage agent’s behavioral shifts from model migration. Organizations often run legacy models alongside newer versions because agent scaffolds and evaluation suites depend on the specific behaviors of the older model, where sudden updates might degrade output quality. Additionally, governance policies enforce teams to route subtasks to different model endpoints based on user or developer access levels. Thus, architectural complexity often reflects strategic operations rather than task difficulty.
Interestingly, we observe a heavier tail towards agents using more distinct models when we examine the full survey data, including prototyping and research agents that have not yet been deployed (Figure 16(a)). Deployed agents are more likely to converge on fewer number of distinct models compared to non-deployed agents, suggesting that teams explore richer multi-model combinations during early experimentation but consolidate onto a smaller set of models as they move toward deployment. We hypothesize this might be also reflecting the additional operational burden of maintaining many distinct model endpoints.
5.2 Model Weights Tuning
@@
We observe a strong preference for prompting over model weight updates in deployed agents. We find that 14 out of 20 (70%) detailed case studies rely solely on off-the-shelf models without supervised fine-tuning (SFT) or reinforcement learning (RL) (Figure 6). Additionally, 2 teams from detailed case studies explicitly report that foundation model capabilities already meet their target use case, making fine-tuning unnecessary.
Only 5 of 20 detailed case studies actively use SFT. These teams target deployment in business-specific corporate contexts where leveraging highly contextual information improves downstream performance. For example, customer product support agents benefit from fine-tuning on specific product offerings and policies. However, performance gains do not always justify the development overhead. Among the 5 detailed case studies actively using fine-tuned models, 3 consider SFT essential, while 2 apply it selectively for enterprise clients where customization requirements justify the additional cost. Three additional teams from interview case studies express strong interest in future adoption for similar reasons. Notably, we find that 4 of the 5 detailed case studies that employ SFT do so in combination with off-the-shelf LLMs, rather than relying on fine-tuned models exclusively.
Regarding reinforcement learning (RL), only 1 scientific discovery case from our interviews uses an RL post-trained model. Three other teams express interest in exploring RL for software testing in future development cycles.
This data, however, does not diminish the value of post-training models for agent applications. Interviews show that SFT and RL is challenging to implement and brittle to model upgrades. Given that off-the-shelf base models can already do most of what the agent applications need, teams prefer methods with lower integration overhead that do not increase already high development and maintenance burdens.
5.3 Prompting Strategies
We find that humans dominate system-prompt construction in production systems. Our survey data reveals that 33.9% of deployed agents use fully manual methods with hard-coded strings. Another 44.6% use a hybrid approach where humans manually draft prompts and then use an LLM to augment or refine them, and 3.6% rely on utilizing predefined prompt templates. Only 8.9% of respondents use a prompt optimizer (e.g., DSPy [khattab2024dspy]) to improve their agent systems, and just 3.6% report letting agents autonomously generate their own prompts.
Our detailed case studies confirm this pattern. Only 1 out of 20 (5%) detailed case studies has explored automated prompt optimization. The remaining cases rely on primary human construction, sometime using LLMs for augmentation. While recent research [zhang2025agenticcontextengineeringevolving, agrawal2025gepareflectivepromptevolution, khattab2024dspy] proposes automating prompts into parametric optimizations, we find these methods rare in deployment. We speculate from interview conversations with practitioners that they prioritize controllable, interpretable methods that allow for fast iteration and debugging over automated or “black-box" methods that requires additional engineering overhead.
We find that prompt complexity increases with system maturity. Among deployed agents from survey, we observe a wide distribution: while 51.5% of systems use short instructions under 500 tokens, there is a long tail of massive prompts (Figure 8(b)). Specifically, 24.2% of deployed agents use prompts between 500 and 2,500 tokens, 12.1% use between 2,500 to 10,000. Notably, another 12.1% even exceed 10,000 tokens .
@@ 5.4 Agent Architecture
We explore the core architectural patterns that support production deployment. To ensure clarity, we adopt the terminologies visualized in Figure 22: an agent completes a high-level task, which decomposes into logical subtasks, consisting of granular atomic steps (e.g., model calls, tool use).
Number of Steps. We find that production agents tend to follow structured workflows with bounded autonomy. We ask survey participants how many steps their deployed systems execute within a subtask before requiring human input. Most systems operate within tight bounds: 46.7% of deployed survey agents complete only 1–4 steps, and an additional 21.7% perform 5–10 steps (Figure 8(c)). A smaller subset (16.7%) extends to tens of steps, while only 6.7% report systems with no explicit step limit. Interestingly when we expand the analysis to include all agents including both deployed and not yet deployed agents in Figure 16(c), the distribution shifts toward substantially higher step counts. This indicates that prototypes and research systems are more likely to run tens of steps or have no explicit limit on autonomous cycles, reflecting more aggressive exploration of open-ended autonomy during early development.
Practical constraints drive this design choice. Case study participants identify problem complexity, non-determinism in agent self-planning, and latency requirements as key limiting factors. Practitioners intentionally impose limits on reasoning steps to maintain reliability and manage computational time and costs. This simplicity reflects a broader preference for predictable, controllable workflows over experimental open-ended autonomy in production environment.
Number of Model Calls. While distinct from logical steps (which often include non-inference actions like tool execution), we specifically analyze model calls to gauge the inference intensity of deployment systems. We observe that within a single subtask, deployment systems typically execute model calls on the order of tens or less. The majority (66.7%) of deployed survey agents use fewer than 10 calls per subtask, with 46.7% using fewer than 5 calls. This is followed by 33.3% using tens of calls, 9.0% in the hundreds, and 6.1% in the thousands. Interestingly, another 12.1% do not set explicit limits. This aligns with our detailed case studies, where 3 teams confirm executing tens of model calls per subtask (up to 40–50 calls).
Experimental systems are far more likely to sit in the long tail, with many agents routinely making tens of calls. In contrast, deployed agents concentrate in the lower-call regime, suggesting that teams aggressively cap or refactor call budgets as they move from experimentation to deployment to probably control cost, latency, and failure amplification.
Despite the pattern of limited model calls, 31% of deployed survey agents already use various inference-time scaling techniques, compared to 44% in experimental systems. While this figure currently includes simpler methods like composing outputs from multiple models (Section 5.1), future work may determine how advanced techniques—such as self-planning, search-based reasoning, and self-verification—perform in production, as these approaches may then lead to higher numbers of model calls and steps before human intervention .
Agent Control Flow. We observe that production agents favor predefined static workflows over full open-ended autonomy. We find that 80% of our detailed case studies utilize a structured control flow. These agents operate within well-scoped action spaces rather than freely exploring the environment to self-determine objectives.
Detailed case studies provide insight into these control flow patterns. Nine cases utilize various forms of agentic Retrieval-Augmented Generation (RAG) pipelines, ranging from single agents retrieving information via tool calls to sophisticated pipelines with over 20 subtasks that explicitly configure retrieval at certain steps. For example, one insurance agent follows a fixed sequence: coverage lookup, medical necessity review, and risk identification.‡‡‡This is a simplified example workflow to illustrate the core logic, redacted to protect the anonymity of the interviewee. While the agent possesses autonomy to complete each subtask (e.g., deciding if a case needs human intervention for risk identification), the high-level objective and expected output of each subtask remain fixed.
Open-ended autonomy remains rare. We observe only one case where agents operate with unconstrained exploration. Notably, this system runs exclusively in a sandbox environment with rigorous CI/CD verification on the final outputs, avoiding direct interaction with production environment.
However, we identify a growing interest in expanding autonomy. Four detailed case studies employ a planning and routing agent to decompose input requests and dispatch them to task-specialized agents. Another team specializes agents into generators and verifiers, enabling greater autonomy through automated verification. Several teams share that they are experimenting with flexible workflows by allowing agents to make autonomous decisions about next steps or by using planning and orchestration agents.
5.5 Agentic Frameworks
We find a divergence in framework adoption between survey respondents and interview case studies. Among deployed agents from the survey, two-thirds (60.7%) use third-party agentic frameworks. Reliance concentrates around three primary frameworks: LangChain/LangGraph [langchain, LangChain2025LangGraph] leads with 25.0%, followed by CrewAI [CrewAI2025Homepage] at 10.7%, with LLaMAIndex [LlamaIndex2025Homepage] and OpenAI Swarm [OpenAI2025SwarmGitHub] both at 3.6% (Figure 9).
In sharp contrast, our detailed case studies reveal a strong preference for custom in-house agent implementations. Only 3 of 20 (15%) detailed case studies rely on external agent frameworks (2 use LangChain, 1 uses BeeAI). The remaining 17 teams (85%) build their agent application entirely in-house with direct model API calls. For example, one interview case explicitly shared that their agents are their own implementation of ReAct loops. Notably, two additional teams report starting with frameworks like CrewAI during the experimental prototyping phase but migrating to custom in-house solutions for production deployment to reduce dependency overhead.
6 3 Results: How Are Agents Evaluated For Deployment?
Evaluation practices shape which agentic systems reach production and how teams iterate on deployed systems. We investigate how practitioners evaluate and test their agents to meet production deployment requirements, examining both offline evaluation during development and online evaluation in production environments. Specifically, we examine two aspects: what practitioners compare their systems against (baselines and benchmarks), and what methods they use to verify system outputs (evaluation methods).
Our findings reveal that evaluation practices vary widely across production agents, even within the same application domain, shaped by the specific requirements of each deployment context and the availability of ground truth data. Notably, practitioners currently focus on agent output quality and correctness rather than traditional software reliability metrics. Based on 20 detailed case studies, no team reports applying standard production reliability metrics such as five 9s availability to their agent systems. Instead, evaluation centers on whether agents produce correct, high-quality responses.
@@ 6.1 Baselines and Benchmarks
During development, teams conduct offline evaluation to assess agent performance before deployment. Figure 10(a) shows that 38.7% of survey respondents compare their deployed agentic systems against non-agentic baselines such as existing software systems, traditional approaches, or human execution. The remaining 61.3% do not perform baseline comparisons. Among these, 25.8% report their systems are truly novel with no meaningful baseline for comparison. While we do not know reasons behind why the remaining 35.5% did not conduct the comparison, in-depth interviews reveal that baseline comparison is often challenging even when alternatives exist. One reason is that non-agentic baselines frequently combine multiple components, making systematic technical comparison difficult. The HR support agent development team illustrates that agent solution replace a baseline process combining company document lookup, human procedures, and non-LLM HR software. While outcomes such as task completion time are measurable, isolating technical performance for comparison is challenging.
Beyond baseline comparison, teams employ benchmark evaluations. We refer to benchmark here as a curated set of tasks or questions with known correct answers. Among 20 teams, 5 build custom benchmarks. One team builds benchmarks from scratch, collecting ground truth labels through collaboration with domain experts. Four teams synthesize existing data to curate benchmarks, drawing from past test cases, system logs, pull requests, and support tickets. One team also leverages public benchmarks during early development. The remaining 15 teams (75%) evaluate their agents without benchmark sets, using alternative methods such as A/B testing, user feedback, and production monitoring.
Among the 5 teams building custom benchmarks, our interviews reveal a prevalent evaluation pattern despite diverse domains and organizations (e.g., human resources, cloud infrastructure, and business analytics). Teams establish a golden question-and-answer set, collect user interaction and feedback data, then work with subject matter experts to examine quality and expand the golden set. This process extends into production runtime, enabling LLM-as-a-judge online evaluation pipelines. Based on our analysis, the convergence of nearly identical pipelines across diverse contexts suggests research and development opportunities in reusable data ingestion pipelines, curation methods for golden sets, and synthetic generation techniques for evaluation datasets.

### v2

@@ 4.2 Application Domains
Among 69 surveyed deployed agents, we observe 26 distinct application domains, extending well beyond software engineering. Figure 2 shows Finance & Banking (44%), Technology (48%), and Corporate Services (42%) lead, signaling early adoption. A substantial long tail follows across retail, healthcare and other service industries. This growing diversity suggests that many real-world tasks beyond traditional benchmarks (e.g., coding, mathematical reasoning) are viable candidates for agent applications. We believe the breadth of production agents signals opportunities and demand for research to advance agents for diverse contexts.
@@ 4.3 Users of Agents
The deployed agents surveyed primarily serve human users: 92.5% target humans rather than other software systems (Figure 8). Internal employees comprise 52% of users, external customers 40%, and non-human systems only 8%. Case study interviews reveal this human-centric design may be deliberate. Interviewed practitioners report deploying internally first to mitigate reliability and security risks, where agent errors have lower consequences and human oversight is more available. Even external-facing systems typically augment domain experts rather than replacing human workers, with humans serving as final verifiers of agent outputs. Among surveyed systems, deployment scale varies: 43% serve hundreds of users while 26% serve tens of thousands to over 1 million daily users (Figure 14(c)).
@@ 4.4 Latency Requirements
Survey data shows production agents tolerate surprisingly relaxed latency: 66% allow response times of minutes or longer, and 17% set no explicit limit (Figure 3). This pattern challenges mainstream optimization goals in machine learning systems research focused on latency reduction.
Interview data reveals the tolerance stems from a dominant use case: background automation of human workflows. Interviewed practitioners report that their minutes-scale agents still outperform human baselines by 10x, which is critical when staffing shortages exist and the automated tasks are secondary to human users’ core job responsibilities. For example, C01 deploys agents to automate clinicians’ preparation and obtaining of insurance approval, C02 assists sales personnel with customer care, and C16 helps software engineers triage incidents. Fifteen of 20 case studies can operate asynchronously; some even batch process requests hourly or overnight. For these applications, minute-scale latency beats the alternative human completion time.
Only 5 of 20 cases require real-time responsiveness. These include voice agents operating at human conversation speeds (C04-05), where latency becomes the primary deployment challenge (Section B.4.2). For the majority, relaxed latency requirements allow practitioners to prioritize output quality and reliability over latency optimization. This pattern suggests opportunities for agents and techniques that trade speed for correctness and downstream performance.
5 RQ2 Results: What Models, Architectures, And Techniques Are Used?
We now examine five critical design decisions: model selection, model weights tuning methods, prompt construction, agent architectures, and development frameworks.
@@ 5.1 Model Selection
Case study data shows deployed agents rely heavily on proprietary frontier models. Seventeen of 20 case studies use closed-source models (Figure 4), with 10 teams explicitly reporting Anthropic Claude (Sonnet 4, Opus 4.1) or OpenAI GPT (o3), the state-of-the-art at interview time. Open-source adoption (3/20 cases) addresses specific constraints, such as high-volume workloads where inference costs are prohibitive (C09 fine-tuned a model for infrastructure maintenance using existing company GPUs) or regulatory requirements preventing sensitive data sharing.
Model selection is based on empirical testing. Engineers evaluate the most powerful accessible models and selecting based on downstream performance. These teams report that runtime costs remain negligible compared to alternative expert labor costs (e.g., medical professionals, senior engineers), justifying the use of expensive frontier models.
Number of Distinct Models. Survey data shows that while 41% of deployed agents use a single model, 59% coordinate multiple models. Case study interviews reveal that half of the teams (10/20) combine models driven by both functional needs and surprisingly operational constraints.
Functional needs include (1) cost optimization, routing simple subtasks to smaller models and complex reasoning to powerful models (C16), and (2) modality requirements, combining text-to-speech with LLMs (C04-C05) or pairing domain-specific models (chemistry) with general reasoning models (C06-C08, Figure 13).
Operational constraints. Interview data reveals some teams maintain multiple models to manage fragility from model upgrades. Agent scaffolds, prompts, and evaluations lock onto specific model behaviors. Deploying newer models can break agent workflows, forcing teams to run legacy models alongside updates (C10). This pattern reveals a reliability challenge: in production settings, newer or more capable models do not guarantee improved agent performance.
5.2 Model Weight Tuning
@@
Case study data shows deployed agents overwhelmingly favor prompting over weight tuning. Fourteen of 20 cases (70%) use off-the-shelf models without supervised fine-tuning (SFT) or reinforcement learning (RL) (Figure 4). Two teams explicitly report foundation models already meet their task requirements, making post-training unnecessary.
Five of 20 cases use SFT (Figure 4): two teams apply SFT consistently, targeting business-specific contexts where domain knowledge improves performance (e.g., C17), while three apply it selectively for enterprise clients where training data is available and cost trade-offs make sense (e.g., C14). Even among these five cases, four combine fine-tuned models with off-the-shelf LLMs rather than relying on fine-tuning alone. Only one case uses RL-trained models (C06 for scientific discovery). Three teams report plans for future RL adoption in complex multi-system environments.
Interview data reveals practical considerations that favor prompting over post-training. Teams report that SFT and RL require substantial implementation effort and are brittle to model upgrades, requiring costly retraining when versions change. Interviewed teams prefer methods that reduce development and maintenance overhead.
5.3 Prompt Tuning Strategies
Survey data shows humans remain central to prompt construction despite emerging automated methods. Among surveyed deployed agents, 34% use manual hard-coded prompts, and 45% use manual drafting augmented by LLMs (Figure 5). Only 9% use prompt optimizers (e.g., DSPy (Khattab et al., 2024)). Case study interviews confirm this pattern: only 1 of 20 teams explored automated optimization, while the rest rely on human construction, sometimes with LLM refinement. Interviewed practitioners report prioritizing controllable and interpretable methods enabling fast iteration, whereas we speculate black-box optimizations may incur additional engineering overhead.
Survey data shows prompt complexity varies widely with system maturity. While most deployed agents (52%) use short prompts under 500 tokens, some use very long prompts exceeding 10,000 tokens (Figure 9(b)). Case study interviews reveal that long prompts typically occur in external client-facing agents requiring extensive guardrails.
@@ 5.4 Agent Architecture
Case study data shows predefined structured workflows dominate production deployments. Eighty percent (16/20) of case studies use structured workflows rather than open-ended autonomous planning. These agents operate within scoped action spaces where practitioners define task sequences upfront. Nine cases implement sophisticated agentic RAG pipelines, single agents retrieving via tool calls, or pipelines with 20+ subtasks explicitly configuring retrieval at each step. For example, C01 follows a fixed sequence of coverage lookup, medical necessity review, and risk identification, where the agent autonomously completes each subtask but high-level objectives remain fixed. This pattern reflects practitioners adapting existing business processes into agent workflow rather than building fully autonomous AI workers. Interview data reveal that workflows dominate production to prioritize controllability and human-expert-in-the-loop oversight for reliability. Teams deliberately constrain autonomy for production stability. Only 1 case uses unconstrained exploration, exclusively in sandboxed environments with rigorous CI/CD verification (C09).
Number of Steps Before Human Intervention. Survey data quantifies this constrained autonomy: 68% of deployed agents execute fewer than 10 steps before requiring human intervention, and 47% execute fewer than 5 steps (Figure 6(a)). For comparison, research prototypes show substantially higher step counts (Figure 21(a)), reflecting aggressive autonomy exploration that consolidates during deployment. Interview data shows problem complexity, planning non-determinism, and latency drive intentional step limits.
5.5 Agent Frameworks
As agent systems mature and scale, teams prefer custom agent implementations over third-party frameworks. Among case studies with production deployments, 85% (17/20) build custom in-house implementations with direct API calls; only 3 use external frameworks (LangChain/LangGraph, DSPy). Survey data shows 61% currently use frameworks (led by LangChain/LangGraph at 25%), but 2 teams explicitly migrated from frameworks (e.g., CrewAI) to custom solutions for production deployment.
Why custom over frameworks? Three drivers emerge. First, flexibility: production agents require vertical integration with proprietary infrastructure and data pipelines that rigid frameworks cannot support (C14: bespoke orchestration for varied client environments). Second, simplicity: core agent loops are straightforward to implement directly, avoiding dependency bloat (C12: custom ReAct implementation). Third, security: enterprise policies prohibit certain external libraries, forcing compliant internal solutions.
6 RQ3 Results: How Are Agents Evaluated For Deployment?
We examine how practitioners assess agents for production readiness and runtime correctness, including what teams compare against and verification methods.
@@ 6.1 Baselines and Benchmarks
Survey data shows most teams lack standardized evaluation frameworks during agents development. Only 39% compare their agents systems to non-agentic baselines (existing software, human execution), while 26% report no meaningful baseline exists (Figure 10(a)). Interview data reveals that even when baselines exist, comparison proves challenging, where non-agentic systems often combine multiple components (human procedures, legacy software, document systems), making isolated technical comparison difficult (e.g., C17: HR agent replacing hybrid human-software workflow).
Case study data shows formal benchmarks are even rarer: 75% evaluate without benchmark sets, relying instead on A/B testing, user feedback, and production monitoring. We refer to benchmarks as curated task sets with expected outputs for systematic evaluation. Five teams build custom benchmarks, one from scratch with expert-labeled ground truth, four by synthesizing existing data (test cases, logs, support tickets). These benchmarks are not easily reusable across deployments due to domain-specific tasks. Case studies reveal a convergent pattern across diverse domains (HR, infrastructure, analytics): establish golden question-answer sets, collect user interactions, and iteratively expand with expert review. This pattern extends to agent runtime through LLM-as-a-judge (§6.2), suggesting opportunities for reusable curation methods and synthetic data generation.